OpenStudy (anonymous):

What is the exact area of a circle with diameter 1 mile? (pi)/4 square miles (pi)/2 square miles (pi) square miles 2(pi) square miles

OpenStudy (anonymous):

(pi) = \[\pi\]

OpenStudy (anonymous):

1. \[\pi/4 \] 2. \[\pi/2 \] 3. \[2\pi\] and \[\pi\]